# Enhanced posterior sampling via diffusion models for efficient metasurfaces inverse design

### v2

@@ 2 Physical System
To compute the electromagnetic field interaction with the metasurface, we first obtain the near field using the Finite Difference Time Domain (FDTD) method [8]. The near field is then propagated to the far field using the Fraunhofer diffraction approximation, which applies a Fourier transform to the near field data [24, 25].
The metasurface architecture is engineered to modulate both near- and far-field electromagnetic interactions. Through precise manipulation of the geometry, dimensions, and spatial configuration of subwavelength constituents, tailored electromagnetic responses are realized. This work specifically investigates the optimization of far-field power distribution.
@@ 3 Related works
In this work, we opted to utilize a surrogate network due to its computational speed, differentiability, and scalability properties that are significantly constrained in rigorous methods such as FDTD or RCWA. However, this approach introduces a dependency on the surrogate’s precision, necessitating the development of a highly performant surrogate model as a critical component of this study [7].
These methods are applied to the inverse design of a beam-shaping metasurface composed of pillars with varying radii, targeting a specified far-field amplitude distribution. Optimization results are validated by comparing the simulated far-field response to the target using the $R^{2}$ metric. Given the nonlinear nature of the problem, the $R^{2}$ metric ranges over $(-\infty,1]$, where values closer to 1 indicate better agreement. The evaluation workflow is shown in Figure 1.
@@ 4 Reference methods
In the introduction, several generative methods such as (VAEs) [15] and (GANs) [16]- were discussed. However, these approaches are excluded from comparative analysis due to their demonstrated inadequacy in far-field metasurface inverse design, as evidenced in [34]. Instead, this study evaluates a diffusion model incorporating enhanced posterior sampling and training techniques, benchmarked against a phase retrieval and look-up-table method, as well as gradient descent utilizing a surrogate network.
@@ 6.1 Noising Schedule
As demonstrated by [53], the choice of noise schedule significantly impacts the performance of diffusion models (DMs) across different applications. To optimize far-field precision, we introduced three distinct schedules for the noise variance $\beta(t)$: Linear, Quadratic, and Sigmoid. The function $\beta(t)$ is monotonically increasing, satisfying $\beta(0)=\beta_{\text{start}}$ and $\beta(T)=\beta_{\text{end}}$. Detailed definitions of these schedules are provided in Appendix A.
@@ 6.2 Consistency Loss
Incorporating a consistency loss during the training of diffusion models (DMs) can significantly enhance the quality and reliability of the generated data. DMs, which transform data from a simple prior distribution to a complex target distribution, benefit from additional constraints that ensure coherence and consistency throughout the diffusion process.
@@ 7.1 Posterior Sampling
With $q_{t}=\frac{1}{\sqrt{1-\alpha_{t}}}$. [29] argues that experimental results exhibit greater stability when using the weighting $q_{t}=\frac{\text{cte}}{\|c-S_{\phi}(x_{0|t})\|_{2}}$. This observation is supported by [57], which demonstrates through a Gaussian example that the guidance term tends to be overestimated at the start of the generation when $\sigma_{t}$ is large, and underestimated near the end when $\sigma_{t}$ is small. However, as demonstrated in Figure 3, normalizing the gradient term slightly reduces precision on the final metric for $23\times 23$ pillar metasurfaces. Conversely, when scaling to larger metasurfaces and increasing the dimensionality of the optimization problem, gradient normalization improves performance, as evidenced in Section 9. Posterior sampling yields promising results, as shown in Figure 3. The shape specified as the condition is now recognizable in the verification simulation. Ancestral sampling as proposed in [42] failed to generate shapes that closely resemble the conditioning input. Furthermore, it is possible to combine the consistency loss approach with posterior sampling, as discussed in the section on comparative results.
@@ 7.2 Monte Carlo Posterior Sampling
Monte Carlo posterior sampling [57] is a variation of the posterior sampling method introduced in the previous section. Instead of relying solely on $x_{0|t}$ to compute the guidance term $\nabla_{x_{t}}\|c-S_{\phi}(x_{0|t})\|_{2}^{2}$, $x_{0|t}$ is treated as the mean of a Gaussian distribution, $x_{i,0|t}\sim\mathcal{N}(x_{0|t},r_{t})$, where the hyperparameter $r_{t}$ represents the guidance schedule. This method aims to mitigate the guidance value overshoot and undershoot discussed in the previous section. The new guidance term is then expressed as:
@@
Figure 4: Amplitude of the far field (||Far field||) obtained from inverse-designed metasurface parameters using Monte Carlo posterior sampling with $1,000$ steps and $N=10$. The normalized guidance is defined as $q_{t}=\frac{1}{\|c-S_{\phi}(x_{0|t})\|_{2}}$, while non-normalized guidance corresponds to $q_{t}=1$.
@@ 7.3 Spherical Gaussian Constrained Posterior Sampling
The Spherical Gaussian Constrained Posterior Sampling (PS SG) technique addresses the limitations of previous methods, namely standard posterior sampling (PS) and Monte Carlo posterior sampling (PS MC). Both earlier methods rely on the strong assumption that the data manifold $\mathcal{M}_{0}$ is linear and that the Jensen gap is negligible. However, as demonstrated in [30], this assumption is flawed, with the Jensen approximation introducing errors in the guidance term. The study in [30] highlights that these errors are strongly dependent on the dimensionality of the problem, as indicated by a lower bound.
@@ 8 Comparative Results
Gradient Descent 0.814 0.053
Qualitatively, Figure 5 illustrates the superior performance of the PS SG method over other PS techniques. This observation is quantitatively supported by Table 2, where PS SG demonstrates both the highest average performance relative to reference methods and the greatest consistency, as indicated by its lower standard deviation in the $R^{2}$ metric. For further comparison, the far-field results of GD and Phase Retrieval & Look-Up-Table are provided in Appendix C.
@@ 9 Scaling up
Due to memory and computational speed constraints of FDTD rigorous simulations, the database was limited to small metasurfaces of size 23x23 pillars, significantly smaller than the target size of approximately 100x100 pillars. This choice enabled the generation of a database containing 5k elements within two weeks of simulations, despite the ultimate goal of performing inverse design for metasurfaces of around 100x100 pillars. It is important to note that quantifying the precision of metasurfaces exceeding approximately $100\times 100$ pillars presents challenges due to the substantial computational resources required for rigorous FDTD simulations. Consequently, the surrogate solver and diffusion models (DMs) were trained on the 23x23 pillar database.
@@
Figure 9: Comparison of $R^{2}$ values for the far-field magnitude obtained from inverse-designed metasurfaces using GD, Phase Retrieval & Look-Up Table, and DM methods, evaluated as the number of pillars increases. The GD method consistently underperforms, with its precision declining sharply with larger pillar arrays, leading to its exclusion from the plots due to low accuracy.The results highlight the critical role of guidance term normalization: it ensures that PS and PS MC maintain high precision, whereas their $R^{2}$ values degrade significantly without normalization as the number of pillars grows.
@@ 10 Conclusion
Diffusion models employing basic ancestral sampling proved insufficient for far-field metasurface inverse design. However, the incorporation of advanced posterior sampling techniques, including Monte Carlo (MC) and Spherical Gaussian (SG) constrained sampling, yielded state-of-the-art performance ($R^{2}=0.965$). These methods significantly outperformed raw posterior sampling ($R^{2}=0.916$) and alternative approaches such as gradient descent ($R^{2}=0.814$) and phase retrieval with look-up tables ($R^{2}=0.919$) in terms of both accuracy and scalability.
@@ Appendix A Noising schedule
Figure 10: Evaluation of the data coefficient $\alpha(t)$, computed from $\beta(t)$ using (3), for three noise schedules linear, sigmoid, and quadratic each tested with $\beta_{\text{end}}\in\{0.1,0.01\}$ while fixing $\beta_{\text{start}}=10^{-4}$. The y-axis represents the post-generation performance metric $R^{2}_{\text{far field}}$, assessed using the surrogate network $S_{\phi}$, and the x-axis denotes the training metric, which quantifies the diffusion network’s accuracy in predicting the added noise ($\epsilon$).
@@
First, the post-generation surrogate-simulation performance, measured by $R^{2}_{\text{far field}}$, exhibits significant variability. While certain configurations achieve state-of-the-art results ($R^{2}_{S_{\phi}}\approx 0.99$), others perform below established benchmarks ($R^{2}_{S_{\phi}}\approx 0.65$).
@@
Figure 11: Each blue dot represents a diffusion model trained with distinct schedule parameters drawn from the set $\{\text{schedule function}\}\times\{\beta_{\text{end}}\}\times\{\text{consistency loss}\}=\{\text{sigmoid},\text{quadratic},\text{linear}\}\times\{0.1,0.01,0.001\}\times\{\text{False},\text{scheduled},\text{True}\}$. When combined with enhanced sampling techniques, the best training metric does not necessarily correspond to the best final metric.
@@ Appendix C Method comparison
The far-field condition provided as input to both the diffusion model and the surrogate network is represented as a two-value image, as illustrated in Figure 12. The assigned values are determined by the spatial distribution of light concentration, while the total energy remains consistent across both images in Figure 12.
@@ Appendix D Scaling results in image
In Figures 1415, and 16, the degradation of the far-field pattern relative to the target is evident as the metasurface size increases, particularly for the Gradient Descent method. For this method, the target pattern becomes barely discernible for metasurfaces exceeding $78\times 78$ pillars.

### v3

@@ 2 Physical System
To compute the electromagnetic field interaction with the metasurface, we first obtain the near field using the FDTD method [8]. The near field is then propagated to the far field using the Fraunhofer diffraction approximation, which applies a Fourier transform to the near field data [24, 25].
The metasurface architecture is engineered to modulate both near and far field electromagnetic interactions. Through precise manipulation of the geometry, dimensions, and spatial configuration of subwavelength constituents, tailored electromagnetic responses are realized. This work specifically investigates the optimization of far field power distribution.
@@ 3 Related works
In this work, we opted to utilize a surrogate network due to its computational speed, differentiability, and scalability properties that are significantly constrained in rigorous methods such as FDTD or RCWA [31]. However, this approach introduces a dependency on the surrogate’s precision, necessitating the development of a highly performant surrogate model as a critical component of this study [7].
These methods are applied to the inverse design of a beam-shaping metasurface composed of pillars with varying radii, targeting a specified far field amplitude distribution. Optimization results are validated by comparing the simulated far field response to the target using the $R^{2}$ metric. Given the nonlinear nature of the problem, the $R^{2}$ metric ranges over $(-\infty,1]$, where values closer to 1 indicate better agreement. The evaluation workflow is shown in Figure 1.
@@ 4 Reference methods
In the introduction, several generative methods such as VAEs [15] and GANs [16] were discussed. However, these approaches are excluded from comparative analysis due to their demonstrated inadequacy in far field metasurface inverse design, as evidenced in [34]. Instead, this study evaluates a diffusion model incorporating enhanced posterior sampling and training techniques, benchmarked against a phase retrieval and look-up-table method, as well as gradient descent utilizing a surrogate network.
@@ 6.1 Noising Schedule
As demonstrated by [53], the choice of noise schedule significantly impacts the performance of diffusion models across different applications. To optimize far field precision, we introduced three distinct schedules for the noise variance $\beta(t)$: Linear, Quadratic, and Sigmoid. The function $\beta(t)$ is monotonically increasing, satisfying $\beta(0)=\beta_{\text{start}}$ and $\beta(T)=\beta_{\text{end}}$. Detailed definitions of these schedules are provided in Appendix A.
@@ 6.2 Consistency Loss
Incorporating a consistency loss during the training of DMs can significantly enhance the quality and reliability of the generated data. DMs, which transform data from a simple prior distribution to a complex target distribution, benefit from additional constraints that ensure coherence and consistency throughout the diffusion process.
@@ 7.1 Posterior Sampling
With $q_{t}=\frac{1}{\sqrt{1-\alpha_{t}}}$. [29] argues that experimental results exhibit greater stability when using the weighting $q_{t}=\frac{\text{cte}}{\|c-S_{\phi}(x_{0|t})\|_{2}}$. This observation is supported by [57], which demonstrates through a Gaussian example that the guidance term tends to be overestimated at the start of the generation when $\sigma_{t}$ is large, and underestimated near the end when $\sigma_{t}$ is small. However, as demonstrated in Figure 3, normalizing the gradient term slightly reduces precision on the final metric for $23\times 23$ pillar metasurfaces. Conversely, when scaling to larger metasurfaces and increasing the dimensionality of the optimization problem, gradient normalization improves performance, as evidenced in Section 9. Posterior sampling yields promising results, as shown in Figure 3. The shape specified as the condition is now recognizable in the verification simulation. Ancestral sampling as proposed in [42] failed to generate shapes that closely resemble the conditioning input. Furthermore, it is possible to combine the consistency loss approach with posterior sampling, as discussed in Section 8 on comparative results.
@@ 7.2 Monte Carlo Posterior Sampling
Monte Carlo posterior sampling (PS MC) [57] is a variation of the posterior sampling method introduced in the previous section. Instead of relying solely on $x_{0|t}$ to compute the guidance term $\nabla_{x_{t}}\|c-S_{\phi}(x_{0|t})\|_{2}^{2}$, $x_{0|t}$ is treated as the mean of a Gaussian distribution, $x_{i,0|t}\sim\mathcal{N}(x_{0|t},r_{t})$, where the hyperparameter $r_{t}$ represents the guidance schedule. This method aims to mitigate the guidance value overshoot and undershoot discussed in the previous section. The new guidance term is then expressed as:
@@
Figure 4: ||Far field|| obtained from inverse-designed metasurface parameters using Monte Carlo posterior sampling with $1,000$ steps and $N=5$. The normalized guidance is defined as $q_{t}=\frac{1}{\|c-S_{\phi}(x_{0|t})\|_{2}}$, while non-normalized guidance corresponds to $q_{t}=1$.
@@ 7.3 Spherical Gaussian Constrained Posterior Sampling
The Spherical Gaussian Constrained Posterior Sampling (PS SG) technique addresses the limitations of previous methods, namely standard PS and PS MC. Both earlier methods rely on the strong assumption that the data manifold $\mathcal{M}_{0}$ is linear and that the Jensen gap is negligible. However, as demonstrated in [30], this assumption is flawed, with the Jensen approximation introducing errors in the guidance term. The study in [30] highlights that these errors are strongly dependent on the dimensionality of the problem, as indicated by a lower bound.
@@ 8 Comparative Results
Gradient Descent 0.697 0.053
Qualitatively, Figure 5 illustrates the superior performance of the PS SG method over other PS techniques. This observation is quantitatively supported by Table 2, where PS SG demonstrates both the highest average performance relative to reference methods and great consistency, as indicated by its low standard deviation in the $R^{2}$ metric. For further comparison, the far fields of GD and Phase Retrieval & Look-Up-Table are provided in Appendix C.
@@ 9 Scaling up
Due to memory and computational speed constraints of FDTD rigorous simulations, the database was limited to small metasurfaces of size $23\times 23$ pillars, significantly smaller than the target size of approximately $100\times 100$ pillars. This choice enabled the generation of a database containing 5k elements within two weeks of simulations, despite the ultimate goal of performing inverse design for metasurfaces of around $100\times 100$ pillars. It is important to note that quantifying the precision of metasurfaces exceeding approximately $100\times 100$ pillars presents challenges due to the substantial computational resources required for rigorous FDTD simulations. Consequently, the surrogate solver and DMs were trained on the $23\times 23$ pillar database.
@@
Figure 9: Comparison of $R^{2}$ values for the far field magnitude obtained from inverse-designed metasurfaces using GD, Phase Retrieval & Look-Up Table, and DM methods, evaluated as the number of pillars increases. The GD method consistently underperforms, with its precision declining sharply with larger pillar arrays, leading to its exclusion from the plots due to low accuracy.The results highlight the critical role of guidance term normalization: it ensures that PS and PS MC maintain high precision, whereas their $R^{2}$ values degrade significantly without normalization as the number of pillars grows.
@@ 10 Conclusion
Diffusion models employing basic ancestral sampling proved insufficient for far field metasurface inverse design. However, the incorporation of advanced posterior sampling techniques, including Monte Carlo and Spherical Gaussian constrained sampling, yielded state-of-the-art performance ($R^{2}=0.962$). These methods significantly outperformed raw posterior sampling ($R^{2}=0.941$) and alternative approaches such as gradient descent ($R^{2}=0.697$) and phase retrieval with look-up tables ($R^{2}=0.913$) in terms of both accuracy and scalability.
@@ Appendix A Noising schedule
Figure 10: Evaluation of the data coefficient $\alpha(t)$, computed from $\beta(t)$ using Eq. (3), for three noise schedules linear, sigmoid, and quadratic each tested with $\beta_{\text{end}}\in\{0.1,0.01\}$ while fixing $\beta_{\text{start}}=10^{-4}$. The y-axis represents the post-generation performance metric $R^{2}_{farfield,S_{\phi}}$, assessed on far field using the surrogate network $S_{\phi}$, and the x-axis denotes the training metric, which quantifies the diffusion network’s accuracy in predicting the added noise ($\epsilon$).
@@
First, the post-generation surrogate-simulation performance, measured by $R^{2}_{farfield,S_{\phi}}$, exhibits significant variability. While certain configurations achieve state-of-the-art results ($R^{2}_{farfield,S_{\phi}}\approx 0.99$), others perform below established benchmarks ($R^{2}_{farfield,S_{\phi}}\approx 0.65$).
@@
Figure 11: Each blue dot represents a diffusion model trained with distinct schedule parameters drawn from the set $\{\text{schedule function}\}\times\{\beta_{\text{end}}\}\times\{\text{consistency loss}\}=\{\text{sigmoid},\text{quadratic},\text{linear}\}\times\{0.1,0.01,0.001\}\times\{\text{no consistency},\text{scheduled consistency},\text{consistency}\}$. When combined with enhanced sampling techniques, the best training metric does not necessarily correspond to the best final metric.
@@ Appendix C Method comparison
The far field condition provided as input to both the diffusion model and the surrogate network is represented as a two-value image, as illustrated in Figure 12. The assigned values are determined by the spatial distribution of light concentration, while the total energy remains consistent across both images in Figure 12.
@@ Appendix D Scaling results in image
In Figures 1415, and 16, the degradation of the far field pattern relative to the target is evident as the metasurface size increases, particularly for the Gradient Descent method. For this method, the target pattern becomes barely discernible for metasurfaces exceeding $78\times 78$ pillars.